# A Comprehensive Survey on Data Augmentation

### v3

@@ IV-A1 Color Perturbation
Pixel Erasing masks out all color information in an area of the data; Photometric Transformation imposes a soft adjustment to change the appearance of the image, but the essential information remains visible.
@@
As illustrated in Fig. 2b, the idea of pixel erasing is to mask some information from the original input space or the feature space. Cutout [9] and random erasing [10] first selects a rectangle region and erases all pixels in that region. The erased region is then filled with a color or Gaussian noise. Random Erasing typically adds one or two big rectangular masks to the image, and some other methods deploy more flexible implementations. Hide-And-Seek [38] randomly removes multiple parts of the image so that the model has to learn from multiple parts of the image other than the most discriminating part. Similarly, the Grid-Mask method [39] generates a set of squared masks in the grid layout. It can prevent excessive deletion and reservation and achieve a reasonable balance between deleted and reserved regional information. In most research, random erasing is used in the spatial domain, while REF [40] extends the use of random erasing to the frequency domain. It uses the discrete Fourier transform to process the image, applies random erasing on the image’s frequency domain, and uses the inverse discrete Fourier transform to recover the image to the spatial domain.
@@
Augmentation that changes the overall color and style of the image is called photometric transformation, as shown in Fig 2c. The simplest way is to adjust the brightness of the image. Color inversion replaces the pixels with their “opposite” color. Color casting applies different operations in different channels, for example, increasing the image’s red channel while keeping the green channel and blue channel untouched. Noise injection adds Gaussian or other types of noises to the images. Greyscaling transforms a colored picture into black-and-white by combining channels for different colors, forcing the model to focus on the shapes or edges of target objects instead of their colors. Histogram equalization [41, 42] enhances the image by enhancing its contrast, and white balancing simulates alternative light cast on the object. Kernel filters are matrices that slide across and are convoluted with the original image to generate new images, either sharpening or blurring the image.
@@ IV-A2 Wording Perturbation
Text data augmentation of this type uses word-level manipulation. It applies changes to the choice of words or adds or removes words. Some examples are presented in Table II.
@@
Token replacement refers to using alternative wording in a given sentence. EDA [11] proposes a set of text augmentations, including synonym replacement. It randomly selects several words from a sentence, substituting them with their synonyms, i.e. words that mean exactly or nearly the same. [43] uses WordNet [44], a lexical database to find synonyms for verbs and nouns in sentences and thus make replacements. It also derives word embeddings by Word2Vec [45] and replaces a word with others that have high cosine similarity. Another possibility is to use a knowledge graph. A knowledge graph describes the relationships of entities, and by utilizing the “is equivalent” relation, we can find a series of synonyms for a given word.
@@
Token addition refers to adding tokens to the original text, either introducing junk tokens as noise or enriching the sentences and making them more informative. The random insertion, proposed in EDA [11], inserts a random synonym of a random word at a random position in the sentence. The words added in the EDA play the role of data noising and may carry either little or harmful information. As a variant of EDA, the technique of AEDA [46] only inserts punctuation marks.
@@
Random deletion [11] randomly removes some words from a sentence. In [47], noising techniques are used to avoid overfitting, and the randomly selected word is replaced by a blank placeholder token “_”, rather than simply being removed. In [48], apart from word dropout which is similar to random deletion, it also uses word blanking to replace a word with a special “¡BLANK¿” placeholder.
IV-A3 Node Attributes Perturbation
Graph nodes often contain rich attributes in the feature matrix. This matrix is generally similar to tabular data, so augmenting node attributes does not need to consider the specialty of graph data.
Attribute Masking and Noise Injection
Various typical augmentation approaches on node attributes from random masking [49, 50, 51] to perturbing attributes with random noises [52] can be applied, as we do for common tabular data. These methods are unified in Eq. 5, where $\mathbf{M}\in\{0,1\}^{N\times N}$ is a binary mask, $\odot$ denotes the Hadamard product, and $\mathbf{N_{W}}\sim\mathcal{N}(0,\Sigma)$ is the Gaussian white noise.
$\mathbf{\tilde{X}}=\mathbf{X}\odot\mathbf{M}+\mathbf{N}_{W}$ (5)
Some other methods are more sophisticated. For example, [53] first learns the topology information of the graph by random walk, and uses the topology embedding together with the original features as inputs to a dual graph neural network. Similarly, LA-GNN [54] applies a local augmentation mechanism. It samples features from nodes’ neighborhoods as additional information when augmenting the attributes. A different approach is proposed in DropMessage [55]. Instead of directly modifying the attributes on the nodes, it applies random dropping in the message-passing phase.
@@ IV-A4 Cell Value Perturbation
Tabular data is much simpler compared with other modalities, so it is hard to design complex augmentation methods that make use of spatial or sequential relationships within a data sample as we do for image or text data, and most data augmentation for tabular data perturbs the value stored in the table cells. Table Masking is a simple technique often applied for tabular data augmentation. These corrupted data are often used with the original data under a self-supervised learning framework to learn stronger representations of the dataset. Another approach to tabular data augmentation borrows ideas from Feature Engineering. Some work on Feature Engineering automates feature selection and transformation to generate optimal feature space, and the operations can also be used to augment tabular datasets.
@@
VIME [56] uses a mask generator to corrupt features. In a self-supervised scenario, two pretext tasks respectively reconstruct the original data and estimate the mask vector. The model updates its weights based on the reconstruction loss and the mask vector estimation loss of the two pretext tasks. VIME also considers a semi-supervised scenario. In that case, multiple masks are generated and applied to the feature of the unlabeled samples. Despite not being labeled, the model predictions on these corrupted samples should be similar, and thus a consistency loss is calculated. For labeled samples, a supervised loss is calculated based on the prediction. The two losses jointly contribute to the back-propagation phase. SCARF [57] applies data augmentation in a contrastive learning framework. Its corruption process is similar to that of VIME but it uses InfoNCE loss to evaluate the similarity between the embeddings of the corrupted and original views and then update the model accordingly. MET [58] works in a pattern similar to VIME and SCARF, which calculates the reconstruction loss of a partially masked sample.
@@
Feature transformation [59, 60] and feature generation techniques can effectively enlarge tabular datasets by generating tables with different feature sets [61]. Powered by reinforcement learning, GRFG [62] transforms existing features with cascading agents. It uses arithmetic operations to provide better traceability and explainability. MOAT [63] extends feature transformation to continuous space. It trains an encoder to project the feature transformation process to continuous embeddings and decodes the optimized embeddings for the best feature transformation strategies. LFG [64] and TIFG [65] leverage LLM agents to generate new features that are both explainable and effective. TFWT [66] computes weights for samples and features, and the weights are combined with the data for later tasks. Although these studies focus on feature engineering, their way of thinking and application of reinforcement learning is very enlightening for data augmentation tasks.
@@ IV-A5 Sequence Value Perturbation
The value information of a sequence is carried on the Y-axis, represented by the amplitude at specific timestamps or frequency channels. Amplitude Perturbation augments time-series data by altering the Y-axis values, and Decomposition decomposes the value of a sequence into different parts that express the sequential values from different aspects.
@@
Jittering adds random noise to the value at each timestamp of a sequence. In [67], jittering is used to simulated additive noise $\epsilon$ of wearable sensors:
$\tilde{x}(\epsilon)=\{x_{1}+\epsilon_{1},\cdots,x_{t}+\epsilon_{t},\cdots,x_{T}+\epsilon_{T}\}$ (6)
Scaling [67] alters the data by multiplying the magnitude of the entire sequence with a scalar $\alpha$.
$\tilde{x}(\alpha)=\{\alpha x_{1},\cdots,\alpha x_{t},\cdots,\alpha x_{T}\}$ (7)
Magnitude warping [67] is similar to scaling, but it uses a smooth curve to warp the sequence so that magnitude at different timestamps is scaled with different parameters.
$\tilde{x}(\alpha)=\{\alpha_{1}x_{1},\cdots,\alpha_{t}x_{t},\cdots,\alpha_{T}x_{T}\}$ (8)
Rotation [67], or flipping, assumes the data is symmetric with respect to the X-axis, and makes the data upside down by applying a rotation matrix $R$.
$\tilde{x}(R)=\{Rx_{1},\cdots,Rx_{t},\cdots,Rx_{T}\}$ (9)
Apart from the amplitude, the phase spectrum provides another option for data augmentation. RobustTAD [68] converts the time-series data from the time domain to the frequency domain, as shown in Eq. 10, and then increases all phases $\theta(\omega_{k})$ by a small perturbation.
$F(\omega_{k})=\frac{1}{N}\sum^{N-1}_{t=0}{x_{t}e^{-j\omega_{k}t}}=A(\omega_{k})exp[j\theta(\omega_{k})]$ (10)
@@
Empirical Mode Decomposition (EMD) [69] is used to produce a set of intrinsic mode functions (IMFs). This decomposition relies on extracting the energy linked to different time scales that are inherent to the data. EMD has been utilized as a method of data augmentation in [70] for the purpose of categorizing impact noise in vehicles. Another way to decompose time-series data is STL [71], which decomposes signals into seasonal, trend, and remainder:
$y_{t}=\tau_{t}+s_{t}+r_{t},\quad t=1,2,\cdots N$ (11)
Other works such as [72] and [68] focus on improving robustness with the help of STL components.
@@ IV-B1 Shape Transformation
This type of data augmentation manipulates the spatial relationship between pixels. Image Cropping breaks the original image into pieces (Fig. 2d), and Geometric Transformation perturbs the structure within the original data (Fig. 2e).
@@
Image cropping selects an area from the original image and removes other parts of the image. It includes manual selection, random cropping, attention-based, aesthetics-based, and change-based approaches [73]. Attention-based approaches use certain attention scores to evaluate a subimage. Regions that contain the main topic or other types of information often have high attention scores. [74] uses the entropy, area ratio, and center distance of the image to evaluate the performance of the image segmentation and accordingly adjust the cropping parameters. [75] uses eye-tracking techniques to capture human fixations and thus infer the location of important contents on the image. Aesthetics-based approaches [76, 77, 78] are similar to quality evaluation. For example, [76] detects several subject regions and one background region, extracts features from the regions, estimates and combines posterior probabilities, and determines the quality score according to these probabilities. A2-RL [79] integrates image cropping into a sequential decision-making process and an aesthetics-aware reinforcement learning model to solve the best cropping problem. Change-based approaches such as [80] take into account the content that is removed and extract two sets of specially designed features to model what is changed after cropping.
@@
As one of the most classical image data augmentation methods, geometric transformation changes the spatial relationship of pixels within an image while keeping the value of each pixel untouched to create more data without extra sampling or labeling. Affine transformations like flipping, shearing, rotation, and translation manipulate the image while preserving the colinearity and ratios of distances. Image flipping reflects the image around its vertical axis, horizontal axis, or both axes, rotation turns the image by an angle about a fixed point. Translation means moving an image in a certain direction. It can introduce more variance in the position of target objects in the image. In non-affine transformation, the spatial relationship between pixels is further changed compared with affine transformations because the colinearity and ratio of distances are broken. Perspective transformation [81] mimics images taken from other views of the object. Image stretching [82] preserves or even enlarges the prominent contents in the image and compresses the background or unimportant foreground information. Distortion [34, 83] also stretches the pixels but often focuses on a certain region. PatchShuffle [84] divides an image into block matrices and shuffles the pixels within each block under a certain probability.
@@ IV-B2 Syntax Transformation
We categorize syntax transformation as: Sentence Cropping that splits a sentence into many pieces, and Sentence Morphing that swaps the location of words or grammatical components. Some examples are presented in Table II.
@@
Sentence cropping [85] divides sentences into several small pieces. It often uses dependency trees to extract the relationships between subjects and objects of the sentence and produces multiple smaller sentences that focus on different parts of the original sentence. This operation breaks the syntactic solidity and semantic comprehensiveness of the entire sentence but preserves local syntactic tags and shallow semantic labels.
@@
These methods manipulate sentence structure by altering the position of words or terms. Random swap [11] randomly chooses two words in a sentence and swaps their position. This perturbs the structure and adds noise to the text data across paragraphs. [86] studies two types of syntactic augmentations: inversion and passivization. Inversion swaps the subject and object in a sentence, and the semantics might be different in most cases. Passivization converts a sentence to its passive form, and the meaning of the new sentence is typically consistent with the original one. Sentence rotation [85] borrows the idea of image rotation, where the image is rotated around its center. In sentence rotation, the root of the sentence is chosen as the center of the sentence, and other sentence fragments are connected to their center via the dependency tree.
TABLE II: Demonstration of some elementary operations for text data augmentation
@@ IV-B3 Topology Transformation
Graph topology structure has immense information representation capacity. Naive Topology Perturbation methods modify a single node or edge component or consider a set of nodes as a whole and modify the substructure. Subgraphing extracts subgraphs from the complete graph. Graph Diffusion and Graph Rewiring are more sophisticated techniques that focus on adjusting the edges to obtain a better graph.
@@
Due to their straightforward idea and easy implementation, node dropping and node addition are among the earliest augmentation approaches in graph learning. As Eq. 12 shows, Node dropping [49] randomly discards some nodes from the graph. Such an operation will also remove the attributes that these nodes carry and the edges on these nodes.
$\tilde{\mathcal{V}}=\text{Drop}(\mathcal{V}),\qquad\tilde{\mathbf{G}}=\{\tilde{\mathbf{A}},\tilde{\mathbf{X}}\}=\{\mathbf{A}[\tilde{\mathcal{V}},\tilde{\mathcal{V}}],\mathbf{X}[\tilde{\mathcal{V}},:]\}$ (12)
As for node addition, an example is MPNN [87]. It introduces a latent master node that is connected to all other nodes, which helps messages pass through the graph network. An edge describes the relationship between two nodes, and by perturbing these relationships through edge dropping and edge addition, we also introduce diversity to the structural information of a graph. DropEdge [88] perturbs the edge set by randomly setting a portion of the adjacency matrix to zero. It first produces a sparse matrix from a random subset of the original edges, then derives the augmented edge set by subtracting the sparse matrix from the original adjacency matrix. Edge masking in [51] constructs a Bernoulli distribution, and randomly masks edges under that probability. FairDrop [89] computes a mask according to the network’s homophily, and uses this mask to guide the dropping process to remove unfair connections. AD-GCL [90] applies a GNN-augmenter to learn how to drop edges so that the mutual information of the augmented graph and that of the original graph is minimized.
@@
Subgraph cropping is based on the idea that graph semantics are partially represented in local structures [49]. Some subgraphing methods, such as those used in GCC [91] and SUGAR [92] use random walk to sample a subgraph from the original graph. GrpahCrop [93] applies a node-centric strategy to crop connectivity-preserved subgraphs and expand the dataset with these augmented subgraphs. These subgraphs are able to maintain some topology characteristics of the original graph. Similarly, SUBG-CON [94] uses the Personalized PageRank algorithm to select the important nodes. NeuralSparse [95] proposes to obtain the subgraph by learning. It uses a sparsification network to generate sparsified subgraphs. Subgraph Substitution perturbs the substructure as a whole of a graph, which is especially beneficial for molecular graphs. A molecule graph represents atoms with nodes and chemical bonds with edges. Functional groups, which often decide the chemical characteristics of molecules, can be represented by substructures. So, by wisely sampling substructures, the global semantics of a molecule graph can be well contained. MoCL [96], for example, utilizes substructure information as one of its augmentation schemes. It performs graph augmentation by replacing functional groups, effectively introducing reasonable noise. GEAR [97] also employs such a technique. It augments molecule graphs by environment replacement, introducing noise from environment subgraphs.
@@
Traditional message passing in graph neural networks only leverages one-hop connections, while the edges of a real graph are often noisy or arbitrarily defined, and many relationships in the networks are indirectly contained in higher-order neighborhoods. To address this issue, GDC [98], as in Eq. $\mathbf{\tilde{A}}=\sum_{k=0}^{\infty}\theta_{k}\mathbf{T}^{k}$ , proposes to calculate a series of transition matrices $\mathbf{T}^{k}$ by means such as random walk and combines these transition matrices with weights $\theta_{k}$ to obtain a new adjacency matrix $\mathbf{\tilde{A}}$ and thus generate a new graph that directly contains some node connection hidden in the original graph. The new graph can also be understood as being processed by a denoising filter because graph diffusion smooths out the neighborhood over the graph [98]. This technique is widely used by self-supervised approaches to transform the graph and generate different views for contrastive learning, such as MVGRL [99], SelfGNN [100], and MERIT [101].
@@
Graph rewiring, as presented in Eq. 13, is initially proposed as randomly swapping edges [102] to improve a network’s robustness and mitigate malicious attacks. Some later research such as smart rewiring [103] takes the degrees of chosen nodes when swapping their edges. Stochastic Discrete Ricci Flow [104] is a curvature-based graph rewiring method, which in each step adds an edge that can improve the most negatively curved edge, and then removes the edge with the most positive curvature. Its results demonstrate the ability of graph rewiring to alleviate the over-squashing problem. DiffWire [105] introduces the CT-LAYER and GAP-LAYER in a GNN model, which respectively learns the commute times and spectral gap of the graph. It rewires the adjacency graph so that the graph’s spectral gap is minimized, and by doing so tries to predict the optimal graph structure for the downstream task. DHGR [106] first learns a similarity matrix from the original graph. Then it rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones. Such an operation can effectively improve the homophily of the graph and thus improve GNN performance. Following DHGR [106], HDHGR [107] computes multiple similarity matrices under different meta-paths and thus generates several rewired meta-path subgraphs.
$\vspace{-2mm}\mathbf{\tilde{A}}=\overbrace{\mathbf{A}\odot(\mathbf{1}-\mathbbm{1}_{r})}^{\text{edge removal}}+\overbrace{(\mathbf{1}-\mathbf{A})\odot\mathbbm{1}_{r}}^{\text{edge addition}}$ (13)
IV-B4 Table Structure Transformation
Table Subsetting

### v4

@@ IV-A1 Color Perturbation
Pixel erasing masks out all color information in an area of the data; photometric transformation imposes a soft adjustment to change the appearance of the image, but the essential information remains visible. Both methods are easy to implement, require none to minimal analysis of the sample, and can be applied to various tasks.
@@
As illustrated in Fig. 2b, pixel erasing masks some information from the original sample. Cutout [10] and random erasing [11] select a rectangular region, erase all pixels in that region, and then fill the region with a color or Gaussian noise. Hide-and-Seek [38] randomly removes multiple parts of the image so that the model has to learn from multiple parts of the image other than the most discriminating part. Grid-Mask [39] generates a set of squared masks in the grid layout. It can prevent excessive deletion and reservation and achieve a reasonable balance between deleted and reserved regional information. REF [40] extends the use of random erasing from the spatial domain to the frequency domain. It uses the discrete Fourier transform to process the image, applies random erasing on the image’s frequency domain, and uses the inverse discrete Fourier transform to recover the image to the spatial domain. Despite its easy application, pixel erasing is subject to removing critical parts from the sample, and some solutions are discussed in Section VII-B1.
@@
As shown in Fig. 2c, photometric transformation changes the overall color and style of the image. The simplest way is to adjust the brightness of the image. Color inversion replaces the pixels with their “opposite” color. Color casting applies different operations in different channels. Noise injection adds Gaussian or other types of noises to the images. Greyscaling transforms a colored picture into black-and-white by combining channels for different colors. Histogram equalization [41] enhances the image by enhancing its contrast, and white balancing simulates alternative light cast on the object. Kernel filters are matrices that slide across and are convoluted with the original image to generate new images, either sharpening or blurring the image.
@@ IV-A2 Wording Perturbation
Token replacement, addition, and deletion perform word-level manipulation. Some examples are presented in Table III.
@@
Token replacement uses alternative wording in a sentence. EDA [12] proposes synonym replacement. It randomly selects words from a sentence, substituting them with their synonyms, i.e., words that mean exactly or nearly the same. [42] uses WordNet [43], a lexical database, to find synonyms for verbs and nouns and make replacements. It also derives word embeddings by Word2Vec [44] and replaces a word with others that have high cosine similarity. KG-NMT [45] resorts to a knowledge graph for terminology equivalence to augment samples and improve translation performance.
@@
Token addition either introduces junk tokens as noise or enriches the sentences to make them more informative. The random insertion, proposed in EDA [12], inserts a random synonym of a random word at a random position in the sentence. The words added in the EDA play the role of data noising and may carry either little or harmful information. As a variant of EDA, the technique of AEDA [46] only inserts punctuation marks.
@@
Random deletion [12] randomly removes some words from a sentence. In [47], noising techniques are used to avoid overfitting, and the randomly selected word is replaced by a blank placeholder token “_”, rather than simply being removed. In [48], apart from word dropout, which is similar to random deletion, it also uses word blanking to replace a word with a special “¡BLANK¿” placeholder.
IV-A3 Node Attributes Masking and Noise Injection
[…]
@@ IV-A4 Cell Value Perturbation
Table masking is a simple technique often applied for tabular data augmentation that corrupts cell values. Another approach borrows ideas from feature engineering, which automates feature selection and transformation to generate optimal feature space, and the operations can also be used to augment tabular datasets.
@@
VIME [56] uses a mask generator to corrupt features. In a self-supervised scenario, two pretext tasks reconstruct the original data and estimate the mask vector, respectively. SCARF [57] applies data augmentation in a contrastive learning framework. Its corruption process is similar to that of VIME, but it uses InfoNCE loss to evaluate the similarity between the embeddings of the corrupted and original views and then update the model accordingly. MET [58] works in a similar pattern, which calculates the reconstruction loss of a partially masked sample.
@@
Feature transformation and generation techniques can effectively enlarge tabular datasets by generating tables with different feature sets [59, 60, 61]. GRFG [62] transforms existing features with cascading agents that search for the best arithmetic operation sets. MOAT [63] extends feature transformation to continuous space by decoding the optimal transformation strategies from continuous embeddings. LFG [64] and TIFG [65] leverage LLM agents to generate new features that are both explainable and effective. TFWT [66] computes weights for samples and features, and the weights are combined with the data for later tasks. Although these studies focus on feature engineering, their way of thinking and application of reinforcement learning are very enlightening for data augmentation tasks.
@@ IV-A5 Sequence Value Perturbation
The value information of a sequence is carried on the Y-axis, represented by the amplitude at specific timestamps. Amplitude perturbation augments time-series data by altering the Y-axis values, and sequence decomposition extracts different parts that express the sequential values from different aspects.
@@
Jittering adds random noise to the value at each timestamp of a sequence. In [67], jittering is used to simulate additive noise $\epsilon$ of wearable sensors: $\tilde{x}(\epsilon)=\{x_{1}+\epsilon_{1},\cdots,x_{t}+\epsilon_{t},\cdots,x_{T}+\epsilon_{T}\}$. Scaling [67] alters the data by multiplying the magnitude of the entire sequence with a scalar $\alpha$: $\tilde{x}(\alpha)=\{\alpha x_{1},\cdots,\alpha x_{t},\cdots,\alpha x_{T}\}$. Magnitude warping [67] resembles scaling but applies a smooth curve to the sequence so that magnitude at each timestamp is scaled with different parameters: $\tilde{x}(\alpha)=\{\alpha_{1}x_{1},\cdots,\alpha_{t}x_{t},\cdots,\alpha_{T}x_{T}\}$. Rotation [67], or flipping, assumes the data is symmetric with respect to the X-axis and makes the data upside down by applying a rotation matrix $R$: $\tilde{x}(R)=\{Rx_{1},\cdots,Rx_{t},\cdots,Rx_{T}\}$. Apart from the amplitude, the phase spectrum provides another option for data augmentation. RobustTAD [68] converts the time-series data from the time domain to the frequency domain, as shown in equation: $F(\omega_{k})=\frac{1}{N}\sum^{N-1}_{t=0}{x_{t}e^{-j\omega_{k}t}}=A(\omega_{k})exp[j\theta(\omega_{k})]$, and then increases all phases $\theta(\omega_{k})$ by a small perturbation.
@@
Empirical Mode Decomposition (EMD) [69] is used to produce a set of intrinsic mode functions. This decomposition relies on extracting the energy linked to different time scales that are inherent to the data. EMD has been utilized as a method of data augmentation in [70] for the purpose of categorizing impact noise in vehicles. Seasonal-trend decomposition (STL) [71] decomposes signals into seasonality, trend, and remainder: $y_{t}=\tau_{t}+s_{t}+r_{t},\quad t=1,2,\cdots N$ [72] and [68] focus on improving robustness with the help of STL components.
@@ IV-B1 Shape Transformation
Image cropping breaks the original image into pieces (Fig. 2d), and geometric transformation perturbs the structure within the original data (Fig. 2e). Both of them manipulate the spatial relationship between pixels.
@@
Image cropping selects an area from the original image and removes other parts of the image. Attention-based approaches use attention scores to evaluate a subimage. Regions with the main topic often have high attention scores. [73] uses the entropy, area ratio, and center distance of the image to evaluate the performance of the image segmentation. [74] uses eye-tracking techniques to capture human fixations and infer the location of important content on the image. Aesthetics-based approaches [75, 76, 77] are similar to quality evaluation. [75] detects several subject regions and one background region, extracts features from the regions, estimates and combines posterior probabilities, and determines the quality score according to these probabilities. A2-RL [78] integrates image cropping into a sequential decision-making process and an aesthetics-aware reinforcement learning model to solve the best cropping problem. Change-based approaches, such as [79], take into account the content that is removed and extract two sets of specially designed features to model what is changed after cropping.
@@
Geometric transformation changes the spatial relationship of pixels within an image while keeping the value of each pixel untouched. Affine transformations like flipping, shearing, rotation, and translation manipulate the image while preserving the colinearity and ratios of distances. Image flipping reflects the image around its vertical axis, horizontal axis, or both axes; rotation turns the image by an angle about a fixed point. Translation means moving an image in a certain direction. It can introduce more variance in the position of target objects in the image. Non-affine transformation allows breaking the colinearity and ratio of distances when reshaping the images. Perspective transformation [80] mimics images taken from other views of the object. Image stretching [81] preserves or even enlarges the prominent contents and compresses the background or unimportant foreground information. Distortion [35, 82] focuses on stretching the pixels in a certain region. PatchShuffle [83] divides an image into block matrices and shuffles the pixels within each block under a certain probability.
@@ IV-B2 Syntax Transformation
We categorize syntax transformation as: sentence cropping that splits a sentence into many pieces and sentence morphing that swaps the location of words or grammatical components. Some examples are presented in Table III.
@@
Sentence cropping [84] divides sentences into small pieces. It uses dependency trees to extract the relationships between subjects and objects of the sentence and produces multiple smaller sentences that focus on different parts of the original sentence. This operation breaks the syntactic solidity and semantic comprehensiveness of the entire sentence but preserves local syntactic tags and shallow semantic labels.
@@
These methods manipulate sentence structure by altering the position of words or terms. Random swap [12] randomly chooses two words in a sentence and swaps their position. [85] studies two types of syntactic augmentations: inversion and passivization. Inversion swaps the subject and object in a sentence, and the semantics might be different in most cases. Passivization converts a sentence to its passive form, and the meaning of the new sentence is typically consistent with the original one. Sentence rotation [84] chooses the root of the sentence’s dependency tree as the center (as in an image) and swaps other fragments.
TABLE III: Demonstration of some elementary operations for text data augmentation
@@ IV-B3 Topology Transformation
Graph topology structure has immense information representation capacity. Naive topology perturbation methods modify a node or an edge. Subgraphing extracts subgraphs from the complete graph. Graph diffusion and graph rewiring are more sophisticated techniques that focus on adjusting the edges to obtain a better graph.
@@
Due to their straightforward idea and easy implementation, node dropping and node addition are among the earliest augmentation approaches in graph learning. Node dropping [49] randomly discards some nodes from the graph: $\tilde{\mathcal{V}}=\text{Drop}(\mathcal{V}),\tilde{\mathbf{G}}=\{\tilde{\mathbf{A}},\tilde{\mathbf{X}}\}=\{\mathbf{A}[\tilde{\mathcal{V}},\tilde{\mathcal{V}}],\mathbf{X}[\tilde{\mathcal{V}}$,:]\}. As for node addition, an example is MPNN [86]. It introduces a latent master node that is connected to all other nodes, which helps messages pass through the graph network. An edge describes the relationship between two nodes, and by perturbing these relationships through edge dropping and edge addition, we also introduce diversity to the structural information of a graph. DropEdge [87] randomly sets a portion of the adjacency matrix to zero. Edge masking in [51] constructs a Bernoulli distribution and randomly masks edges under that probability. FairDrop [88] computes a mask according to the network’s homophily and uses this mask to guide the dropping process to remove unfair connections. AD-GCL [89] applies a GNN augmenter to learn how to drop edges so that the mutual information of the augmented graph and that of the original graph is minimized. Despite its popularity, graph topology perturbation is potentially risky in some cases and must be used carefully. The removal of some crucial edges might influence the connectivity of the graph, thereby limiting the use of some graph algorithms that require the graph to be fully connected. Conversely, adding some inter-class edges might escalate representation collapse caused by the inductive bias of the GNN models.
@@
Subgraph cropping assumes graph semantics are partially represented in local structures [49]. GCC [90] and SUGAR [91] use random walk to sample a subgraph from the original graph. GraphCrop [92] applies a node-centric strategy to crop connectivity-preserved subgraphs and expand the dataset with these augmented subgraphs. SUBG-CON [93] uses the Personalized PageRank algorithm to select the important nodes. NeuralSparse [94] proposes to obtain the subgraph by learning. It uses a sparsification network to generate sparsified subgraphs. Subgraph Substitution perturbs the substructure as a whole of a graph. It is especially beneficial for molecular graphs, which represent atoms with nodes and chemical bonds with edges. Functional groups, often deciding the chemical characteristics of molecules, can be represented by substructures. MoCL [95] utilizes substructure information as one of its augmentation schemes. Its replacement of functional groups effectively introduces reasonable noise. GEAR [96] also employs such a technique. It augments molecule graphs by environment replacement, introducing noise from environment subgraphs.
@@
[…]
@@
[…]
IV-B4 Table Subsetting